The following table shows the numbers of patients classified as underweight, normal weight, overweight and obese according to their diabetes status.

Underweight Normal Weight Overweight Obese
Diabetes 8 34 65 43
No Diabetes 12 85 93 40

1. What proportion of patients is both obese and diabetic?
Answer #1

Solution :

From given table,

Total = 8 + 34 + 65 + 43 + 12 + 85 + 93 + 40 = 380

The proportion of patients is both obese and diabetic is

= 43 / 380

= 0.1132